# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ 5 Performance Analysis of RC-S
For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $min\{n_{h},n_{v}\}\geq 4$, (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ decreases initially and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\textit{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance could be obtained by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
Proofs for Lemma 2 is provided in Appendix B. The condition within the lemma indicates that the conclusion is applicable to an RSS where the length-to-width ratio does not exceed two. This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets under the same area. Lemma 2(i) states that the efficiency of sorting by robots exhibits a trend of first increasing and then decreasing with the increase in the number of workstations. This is because, under the worker load balancing criterion, the average distance from the corner loading stations to each outlet is greater than the stations located in the middle of the four sides. Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape reduce the demand for robots. As aforementioned in section 3, the density of outlets remains constant, thus the distribution of outlets in RSS does not need to be altered as well. Lastly, Lemma 2(iii) provides the range of average distance to be covered in a single delivery within RC-S. This will serve as a foundation for the efficiency analysis in the following sections.
Combining equations (15)-(27), the estimation formula of sorting throughput under given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{occupied}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (28)
If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of system throughput could be calculated by setting the value of $n_{r}$ to infinity and assigning workers to all loading stations:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{occupied}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (29)
We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound in the right side of inequality (29). Here we investigate the properties of this bound.
@@
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has one zero point within the valid range of $n_{l}$. Denote the value as $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing in the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 2 can be derived by incorporating Lemma 1 and Lemma 2, which are depicted in detail in Appendix B. Proposition 2 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and this critical point is not achieved when all loading stations are activated, due to workload balancing and uniformly distributed demands. When the proportion of active loading stations approaches 1, the growth rate of $n^{occupied}_{VP}$ will be slower than the growth rate of the service path length, resulting in a decrease in throughput. We will verify the formula through experiments in the next section.
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S against cooperative A* (CA*) (Silver,, 2021) for traffic management in a warehouse, to better illustrate the superiority of RC-S in terms of AGV service distance and system throughput. CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and based on the order of priorities, algorithm find the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (Phillips and Likhachev,, 2011), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover,other well-known MAPF algorithm frameworks, such as PBS (Ma et al.,, 2019) and RHCR (Li et al.,, 2021), are found less efficient compared to CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and thus are excluded from this comparative analysis.
Table 2: RC-S parameters in simulation
@@
We compare the performance in two scenarios, 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. We activate all possible loading station locations for the experiments, and parcel targets are evenly distributed among outlets. The RC-S parameters, detailed in Table 2, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 50 repeated experiments for each scenario. In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operations. Our algorithms and simulations are coded in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with Intel i7-9700F CPU and 16GB RAM. All processes are run single-threaded.
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
@@
$n_{r}$ Throughput (×$\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
@@
Tables 3 and 4 report the system throughput, average service time and average run-time. The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in larger-scale cases, it is even less than $1/10$ of the latter. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates superior throughput compared to CA*, with the exception observed in a 20+20 aisles scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for environments like RSS, characterized by a compact network and numerous outlets as obstacles. The difference in algorithm performance is primarily due to the following improvements: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, while CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; (2) RC-S limits the total number and maximum density of robots and is free of conflicts, while high robot density around outlets in benchmark algorithm leads to significant resource expenditure on conflict resolution.
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in section 5 across different active ratios of loading stations. We consider 5 scenarios with different network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set as the number of activated VPs in the aisle network, plus 5 per loading station to ensure the stability of operation. Other experimental parameters are specified in section 6.1 in detail.
@@
Figure 15: Robot flow distribution in simulation, $n_{h}=n_{v}=20$
@@
Figure 14 shows the trend of the average service distance as a function of the number of active workstations when $n_{h}=n_{v}=20$. It indicates that the formula and simulation results present the similar trend of first decreasing and then increasing. This suggests that the service frequency of each robot is highest when about half of the loading stations are activated. Figure 14 shows the relative error between formula and simulation results under five scenarios. It can be seen that the error is always within ±$\pm 8\%$, indicating a good fit.
@@
Figure 14 shows the trend of the system throughput as a function of the number of activated workstations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating at the maximum capacity, there is a certain degree of resource waste. Figure 14 shows the relative error between formula and simulation results for five cases, which remains within ±$\pm 15\%$. Figure 15 depicts the average traffic flow distribution during a stable operation of the system in simulation with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We notice that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots in the loading stations in the extreme situation.
6.3 Efficiency underestimating in queueing models
@@
Queueing models have been widely used in analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different work pieces (John A. Buzacott,, 1993). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring work pieces from one machine to another for processing. However, we find that queueing models could not capture the efficiency of robots accurately when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in Zou et al., (2021), where a two-tier RSS model with closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in section 3. Eight scenarios of large-scale network with few robots are tested, using the same experimental setup as in section 6.2. The results are shown in Table 5, where Eq.(28) represents the proposed estimation formula. The CQN model significantly underestimates system throughput in scenarios with few robots. It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it results in unnecessary expenses for surplus robots during off-peak periods.
@@
Throughput (×$\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
@@
we also conduct a experiments on a real case of Deppon Express described in Zou et al., (2021), where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations and 170 robots. The The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ were set according to the data provided by Deppon Express. We compared the performance of RC-S with the results from Zou et al., (2021), as shown in Table 6, where $T_{D}$, $T_{CQN}$ and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under RC-S scheme, respectively. Combined with the experimental results in section 6.1, RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput. Meanwhile, the CQN model effectively captures the congestion in the RSS system implemented by Deppon Express, but underestimates the traffic capacity of the network when the RC-S scheme is applied.
@@ 6.4 Evaluating RC-S across diverse scenarios
In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not spread across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits. The placement of outlets is also restricted by site factors, such as load-bearing columns or the presence of certain equipment. Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across each outlet. In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation. For each set of experiments, we evaluate the overall performance across 50 test runs. Each run includes a 30-minute warm-up period, followed by an observation period where average data is recorded over one hour.
We first compared the impact of different numbers of active loading stations on average throughput. We consider five scenarios with varying network scales: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for sorting process. Other RC-S parameters are specified in detail in Table 2. Figure 16 shows the trend of throughput across the five scenarios. It can be observed that large-scale networks not only have a higher throughput upper bound, but their slope—representing the average efficiency per loading station—is also generally higher. This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
Figure 16: The effect of different numbers of active loading stations
@@
We next evaluate the effect of closing some outlets on system throughput. The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters remaining constant. We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center. To control variates, the number of closed outlets in each scenario is set to 1/9 of the total. Scenarios (2), (3), and (4) correspond to areas 1, 2, and 3 in Figure 17(a), respectively. Results from 50 repeated experiments for each scenario are shown in Figure 17(b). The green dashed line represents the average throughput when all outlets are open. It can be observed that randomly closing outlets has no significant impact on system throughput. Closing outlets in areas 1 and 2 results in a decrease in average throughput, while closing outlets in area 3 leads to a slight increase in average throughput. It is intuitive that closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network. However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources. Since most shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
@@
In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency. In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency for each outlet; however, this cannot entirely eliminate demand heterogeneity. We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution. For outlets with extremely low arrival rates, the situation is similar to the previous experiments involving closed outlets, and their impact can be minimized by random assignment. The two types of outlets are spatially uniformly distributed throughout the network. Keeping other experimental parameters constant, we adjust the proportion of the two types of outlets, and the system performance are shown in Figure 18. The results indicate that with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains efficiency, and the impact of demand heterogeneity on system throughput is negligible. This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow across the network.
Table 7: RC-S performance in large-scale system
@@
[…]

### v4

@@ 5 Performance Analysis of RC-S
For uniformly distributed sorting demands, if the length-to-width ratio is less than 2 and $\min\{n_{h},n_{v}\}\geq 4$, then (i) as $\alpha$ increases, $\bar{l}(n_{h},n_{v},\alpha)$ first decreases and then increases; (ii) when the site size is fixed ($n_{h}\cdot n_{v}=\text{constant}$), a square-shaped site has a smaller average travel distance; (iii) when the site is square-shaped, the range of average travel distance is given by $2D\cdot n_{h}<\bar{l}(n_{h},n_{v},\alpha)<2D\cdot\frac{9}{8}n_{h}$.
The proof of Lemma 2 is provided in Appendix B. The conditions in the lemma indicate that its conclusions is applicable to an RSS where the length-to-width ratio does not exceed two. This is a reasonable setup in practice, as excessively elongated sites accommodate fewer outlets for the same area. Lemma 2(i) states that the sorting efficiency of robots exhibits a trend of first increasing and then decreasing with the increase as the number of loading stations increases. This is because, under the worker load-balancing criterion, the average distance from the corner loading stations to each outlet is greater than that from loading stations located in the middle of the four sides. Lemma 2(ii) suggests that, under the permissible conditions, designing the aisle network in a square shape can reduce the demand for robots. As mentioned in Section 3, the outlet density remains constant, thus the distribution of outlets in RSS does not need to be altered. Lastly, Lemma 2(iii) provides the range of the average distance to be covered in a single delivery within RC-S. This serves as a foundation for the efficiency analysis in the following sections.
Combining Equations (16)–(28), the estimation formula for the sorting throughput under a given configuration $(n_{h},n_{v},n_{w},n_{r})$ is as follows:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})=\frac{D\cdot n^{\text{occupied}}_{VP}(\frac{n_{w}}{n_{h}+n_{v}},n_{r})}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{w}}{n_{h}+n_{v}})}$ (29)
If we fix the values of $n_{h}$ and $n_{v}$, the upper bound of the system throughput can be obtained by setting $n_{r}$ to infinity and assigning workers to all loading stations:
$\tilde{T}_{O}(n_{h},n_{v},n_{w},n_{r})\leq\frac{D\cdot n^{\text{occupied}}_{VP}(\frac{n_{l}}{n_{h}+n_{v}},\infty)}{\tau_{e}\cdot\bar{l}(n_{h},n_{v},\frac{n_{l}}{n_{h}+n_{v}})}$ (30)
We use $\tilde{T}_{M}(n_{h},n_{v},n_{l})$ to denote this upper bound on the right-hand side of Inequality (30). We now investigate the properties of this bound.
@@
$\frac{\partial\tilde{T}_{M}(\cdot)}{\partial n_{l}}$ has a unique zero point within the valid range of $n_{l}$. Denote this value by $n_{l}^{(0)}$, then $\tilde{T}_{M}(\cdot)$ is monotonically increasing on the interval $(0,n_{l}^{(0)}]$ and monotonically decreasing in the interval $(n_{l}^{(0)},n_{h}+n_{v}]$.
The proof of Proposition 3 can be established by combining Lemma 1 and Lemma 2, and is presented in detail in Appendix B. Proposition 3 reveals that under the RC-S robot control scheme, the maximum throughput has a critical point, and that this critical point is not attained when all loading stations are activated, due to workload balancing and uniformly distributed demand. When the proportion of active loading stations approaches 1, the growth rate of $n^{\text{occupied}}_{VP}$ becomes slower than the growth rate of the service path length, resulting in a decrease in throughput. We will verify this formula through experiments in the next section.
@@ 6.1 Comparative analysis: RC-S versus benchmark traffic management framework
In this section, we compare our proposed RC-S with cooperative A* (CA*) (Silver_2021) for traffic management in a warehouse, in order to illustrate the superiority of RC-S in terms of robot travel distance and system throughput. CA* is a framework based on a simple prioritized-planning scheme: each agent is first assigned a unique priority, and, in order of these priorities, the algorithm finds the shortest path for each agent that avoids conflicts with agents of higher priority. We use SIPP (phillips2011sipp), an efficient variant of location-time A*, as the lower-level solver of CA*. Moreover, other well-known MAPF frameworks, such as PBS (ma2019searching) and RHCR (Li_Tinka2021), are found to be less efficient than CA* in large-scale settings ($n_{r}>100$) of our experiments, primarily due to the increased likelihood of cycle conflicts within the RSS framework, and are thus excluded from this comparative analysis.
Table 3: RC-S parameters used in the simulations
@@
We compare the performance in two scenarios: 12+12 aisles and 20+20 aisles. To ensure uniformity in the experimental setup, we use directed maps in all tests to prevent swapping conflicts. All potential loading-station locations are activated in the experiments, and parcel destinations are evenly distributed among outlets. The RC-S parameters, detailed in Table 3, are based on sorting robot data from the Geekplus company. For a fair comparison of algorithm performance, robots move at a constant speed of 2 m/s, matching the speed of VPs in RC-S. The simulations do not consider acceleration and deceleration phases. We conduct 50 independent experiments for each scenario. In each experiment, we set a warm-up period of 30 minutes and then record the system performance for 60 minutes of continuous operation. Our algorithms and simulations are implemented in Python 3.11, and all experiments are conducted on a personal computer running Windows 11 with an Intel i7-9700F CPU and 16 GB of RAM. All processes are run single-threaded.
@@
$n_{r}$ Throughput ($\times 10^{3}$/h) Service time (s) Run-time per cycle(ms)
@@
$n_{r}$ Throughput ($\times 10^{3}$/h) Service time (s) Run-time per cycle (ms)
@@
Tables 4 and 5 report the system throughput, average service time, and average run-time. The comparative results indicate that the run-time of RC-S is consistently lower than that of CA*, and in large-scale cases, it can be less than one tenth of that of CA*. Moreover, RC-S always assigns shorter service paths to robots, resulting in an average service time reduction of 10.3% compared to CA*. In most experiments, RC-S demonstrates higher throughput than CA*, with the only exception observed in the 20+20–aisle scenario with $n_{r}=400$. This observation underscores the suitability of the RC-S algorithm for RSS, which is characterized by a compact network and numerous outlets acting as obstacles. The difference in algorithm performance is primarily due to the following factors: (1) RC-S optimizes by selecting from candidate spatio-temporal paths, whereas CA* and similar cell-based search algorithms consider variable waiting times at each cell and allow robots to return to previously visited cells; and (2) RC-S imposes limits on both the total number and the maximum density of robots and remains conflict-free, whereas the high robot density around outlets in the benchmark algorithm leads to significant resource expenditure on conflict resolution.
@@ 6.2 Throughput estimation formula validation
This section validates the throughput estimation formula proposed in Section 5 for different active ratios of loading stations. We consider five scenarios with different network sizes: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set to be equal to the number of activated VPs in the aisle network, plus five robots per loading station to ensure stable operation. Other experimental parameters are specified in detail in Section 6.1.
@@
Figure 16: Robot flow distribution in simulation, $n_{h}=n_{v}=20,n_{w}=40$
@@
Figure 15 shows the trend of the average service distance as a function of the number of active loading stations when $n_{h}=n_{v}=20$. It indicates that the formula and the simulation results exhibit a similar trend of first decreasing and then increasing. This suggests that the service frequency of each robot is highest when about half of the loading stations are activated. Figure 15 shows the relative error between the formula and the simulation results under five scenarios. It can be seen that the error is always within $\pm 8\%$, indicating a good fit.
@@
Figure 15 shows the trend of the system throughput as a function of the number of activated loading stations when $n_{h}=n_{v}=20$. The trend demonstrates a clear marginal effect of the number of activated loading stations: when the system is operating near its maximum capacity, there is a certain degree of resource waste. Figure 15 shows the relative error between the formula and the simulation results for five cases, which remains within $\pm 15\%$. Figure 16 depicts the average traffic flow distribution during stable operation of the system in simulations with the maximum number of workers. The darker the color of the aisle segments, the higher the traffic load. We observe that the corner loading stations have almost no incoming or outgoing traffic flow. This indicates that Algorithm 1, in practical execution, cannot effectively balance the distribution of robots over loading stations in this extreme situation.
6.3 Underestimation of efficiency in queueing models
@@
Queueing models have been widely used in the analysis of flexible machining systems, where flexibility is defined as the ability to arbitrarily specify the processing sequence of different workpieces (buzacott1993stochastic). In robotic systems, robots function similarly to pallets in traditional manufacturing systems, transferring workpieces from one machine to another for processing. However, we find that queueing models could not accurately capture robot efficiency when the number of robots has not reached the network’s capacity limit. We apply the CQN model established in zou2021robotic, in which a two-tier RSS with a closest Robot-to-Loading-Station assignment rule is modified to meet the assumptions in Section 3. Eight scenarios of large-scale networks with few robots are tested, using the same experimental setup as in Section 6.2. The results are shown in Table 6, where Eq. (29) represents the proposed estimation formula. The CQN model significantly underestimates system throughput in scenarios with few robots. It lacks the capability to accurately depict traffic capacity under low flow conditions with RC-S control and overestimates the impact of conflicts under low-demand conditions, thereby underestimating individual robot efficiency. Although the CQN model provides accurate estimates when the network approaches capacity, it leads to unnecessary costs due to surplus robots during off-peak periods.
@@
Throughput ($\times 10^{3}$/h) Simulation 8.43 16.25 28.60 32.90 7.27 14.30 27.01 36.49
@@
We also conduct an experiment on a real case of Deppon Express described in zou2021robotic, where the network scale is $n_{h}+n_{v}=18+6$, with 108 outlets, 6 loading stations, and 170 robots. The aisle width $D$, maximum robot speed $v_{max}$, and worker loading rate $r_{l}$ are set according to the operational data provided by Deppon Express. We compare the performance of RC-S with the results from zou2021robotic, as shown in Table 7, where $T_{D}$, $T_{CQN}$, and $T_{RCS}$ represent the throughput of the Deppon system, the CQN model and the simulation under the RC-S scheme, respectively. Together with the experimental results in Section 6.1, this real-case study shows that the RC-S scheme significantly improves the operational efficiency of robots, thereby enhancing the overall system throughput. Meanwhile, the CQN model effectively captures congestion in the RSS system implemented by Deppon Express but underestimates the traffic capacity of the network when the RC-S scheme is applied.
@@ 6.4 Evaluating RC-S across diverse scenarios
In real sorting centers, due to cost constraints and demand arrival rates, loading stations are often not distributed across the entire perimeter of the RSS, resulting in only partial availability of network entrances and exits. The placement of outlets is also restricted by site-specific factors, such as load-bearing columns or the presence of certain equipment. Additionally, demand arrival rates are heterogeneous and are not uniformly distributed across outlets. In this section, we evaluate the performance of RC-S across various application scenarios and test the impact of different system configurations on efficiency by simulation. For each set of experiments, we evaluate overall performance across 50 independent runs. Each run includes a 30-minute warm-up period, followed by an observation period during which average performance is recorded over one hour.
We first compare the impact of different numbers of active loading stations on average throughput. We consider five scenarios with varying network sizes: $n_{h}=n_{v}=12,14,16,18,20$, respectively. The number of robots is set to be sufficient in each scenario, ensuring that it does not become a bottleneck for the sorting process. Other RC-S parameters are specified in detail in Table 3. Figure 17 shows the trend of throughput across the five scenarios. It can be observed that large-scale networks not only have a higher upper bound on throughput but also exhibit steeper slopes, representing higher average efficiency per loading station. This is because, under RC-S, larger networks provide more space to distribute incoming traffic across different aisles, thereby reducing congestion at each entrance.
Figure 17: Effect of different numbers of active loading stations on throughput
@@
We next evaluate the effect of closing a subset of outlets on system throughput. The network size is set to $n_{h}=n_{v}=20$, with other experimental parameters kept constant. We consider four scenarios for closing outlets: (1) randomly closing outlets; (2) closing outlets in the corners; (3) closing outlets along the edges; (4) closing outlets in the center. To control for other factors, the number of closed outlets in each scenario is set to one ninth of the total. Scenarios (2), (3), and (4) correspond to Areas 1, 2, and 3 in Figure 18(a), respectively. Results from 50 independent experiments for each scenario are shown in Figure 18(b). The green dashed line represents the average throughput when all outlets are open. It can be observed that randomly closing outlets does not have a significant impact on system throughput. Closing outlets in Areas 1 and 2 results in a decrease in average throughput, while closing outlets in Area 3 leads to a slight increase in average throughput. Intuitively, closing outlets in a specific area causes an uneven distribution of demand, leading to imbalanced network traffic and increased traffic load in other areas of the network. However, from another perspective, it reduces robot turns within the closed area, thereby improving the utilization of spatio-temporal resources. Since most of the shortest paths pass through the central area, closing central outlets in scenario 4 results in higher throughput compared to the normal case.
@@
In addition, we evaluate the effect of heterogeneous demand distribution on system efficiency. In real sorting centers, sorting destinations with higher demand arrival rates are often assigned multiple outlets to minimize differences in the visit frequency between outlets; however, this cannot completely eliminate demand heterogeneity. We assume there are two types of outlets in the RSS: one type has an average demand arrival rate twice that of the other, with both following a normal distribution. For outlets with extremely low arrival rates, the situation is similar to the previous experiments involving closed outlets, and their impact can be minimized by random assignment. The two types of outlets are spatially uniformly distributed throughout the network. Keeping other experimental parameters constant, we adjust the proportion of the outlet types, and the resulting system performance is shown in Figure 19. The results indicate that, with a fixed total demand and varying proportions of high-demand outlets, RC-S maintains its efficiency, and the impact of demand heterogeneity on system throughput is negligible. This is achieved by leveraging demand arrival rates for different destinations and balancing the visit frequency to each area, ensuring a well-distributed traffic flow throughout the network.
Table 8: RC-S performance in large-scale systems
@@
Finally, we evaluate the performance of RC-S in large-scale systems. The experimental scenarios include five large-scale networks with $n_{h}=n_{v}=22,\ 24,\ 26,\ 28$, and $30$, respectively. Among these, the $30+30$ network covers an actual area of over 5,000 square meters, significantly larger than most sorting systems in modern warehouses. For example, a real case from Deppon Express mentioned in the study by zou2021robotic has a network with a scale of $18+6$, approximately one-quarter of the $30+30$ scenario (assuming the same aisle width and loading station space occupation). The results are shown in Table 8, including system throughput, the average and maximum run-time per cycle, and the number of precomputed feasible paths associated with the network size. As the network scale increases, the total number of feasible paths grows, and the path selection time per cycle gradually increases. However, the maximum run-time remains acceptable and is much shorter than the length of each cycle.